Since last Friday either the arms or the PRC can't lock.
The montors show the beam flashing on the end mirrors, but the cavity can't get locked. The error signal looks fine. I suspect a computer problem.
Also PRC can't lock. SPOB is suspiciously stuck at about -95. Although that's not a fixed number, but covering the by hand the SPOB PD on the ITMY table doesn't change the number. I check the DC output of the photodetector and it is actually seen the beam.
Suspecting computer problems started after last Thursday's IP switch, I rebooted the frame builder, c1dcuepics, c1daqctrl and all the front ends. I then burtrestored to February 1st at 1:00 am.
Before I burtrestored c1iscepics, SPOB had gone back to more typical numbers around 0, as it usually read when PRC wasn't locked.
But burtrestoring c1iscepics, return it to the -95 of earlier.
Burterestoring to other times or dates didn't solve the problems.